ROVIBRONIC COUPLINGS IN A FOURFOLD DEGENERATE ELECTRONIC STATE OF AN OCTAHE-DRAL MOLECULE : THE ν_{3} BAND OF ReF_{6}.

\underline{M. REY}, V. BOUDON, M. LOËTE, Laboratoire de Physique de l’Université de Bourgogne, UMR CNRS 5027, 9 Av. A. Savary, BP 4780, F-21078 DIJON Cedex, France.

The rhenium hexafluoride (ReF_{6}) molecule has the particularity to possess an incomplete electronic d subshell. This open-shell system belongs to the O_{h} group and exhibits a very large spin-orbit coupling leading to a fourfold degenerate electronic ground state of symmetry G′_{g}. This implies complex couplings between electronic and nuclear motions (vibronic couplings) and between electronic, nuclear and rotational motions (rovibronic couplings) which considerably complicate calculations. Many years ago, some abnormalities in the spectrum of the triply degenerate ν_{3} band of symmetry F_{1u} (the absence of the usual P QR pattern for such an isolated band, for instance) was attributed to a dynamical quadratic Jahn-Teller effect ^{a}. But all the theoretical vibronic models developed up to now give matrices with very large sizes when adding the rotational part and this leads in practice to inextricable problems. To solve this, we propose a systematic tensorial development in the SU(2) ⊗ C_{I} ⊃ O_{h}^{S} group chain for the construction of the effective rovibronic Hamiltonian^{b}^{,}^{c}. This new theoretical model is based on a polyad scheme and allows a simultaneaous treatment of the four vibronic sublevels

F_{1u} ⊗ G′_{g} = E′_{1u} ⊕ E′_{2u} ⊕ 2G′_{u} .

The corresponding rovibronic energies are obtained by block-diagonalizing the full Hamiltonian for the considered vibronic polyad. The size of each rovibronic block is thus finite. In a previous study, we have successfully tested this model for an electronic triplet state (V(CO)_{6}) ^{d}. Here, we report in a similar way the low resolution band profile of the ν_{3} band of ReF_{6} and show the agreement between the observed ^{e} and calculated spectrum.

aR. S. McDowell and L. B. Asprey J. Mol. Spectrosc. 45, 491-493 (1973).

bM. Rey, V. Boudon, M. Loëte and F. Michelot J. Mol. Spectrosc. 204, 106-119 (2000).

cM. Rey, V. Boudon and M. Loëte J. Mol. Struct. 599, 125-137 (2001).

dM. Rey, V. Boudon, M. Loëte, P. Asselin, P. Soulard and L. Manceron J. Chem. Phys. 114(24), 10773-10779 (2001).

eV. Boudon, M. Rotger, Y. He, U. Schmitt, H, Hollenstein and M. Quack, submitted for publication.